# Online estimation of the inverse of the Hessian for stochastic optimization with application to universal stochastic Newton algorithms

### v1

@@ 6.5 Useful lemmas
Proof of Lemma 6.2.
Note that the case where $E_{n}=0$ is exactly the case of Lemma 6.1. Therefore, we are going to study the case where $E_{n}\neq 0$. We define $\alpha_{n}:=\prod_{k=0}^{n}(1+E_{k})$. Note that since $\sum_{n=0}^{\infty}E_{n}$ converges almost surely, $\alpha_{n}$ converges almost surely to a finite random variable $\alpha_{\infty}$. Moreover, noting
@@
$\mathbb{E}\left[V^{\prime}n|\mathcal{F}_{n}\right]\leq V_{n}^{\prime}+B_{n}^{\prime}-D_{n}^{\prime}.$
@@
According to Lemma 6.1, we have $V^{\prime}_{n}=o(a_{n})$ a.s., and therefore $V_{n}=o(a_{n})$ a.s. Furthermore,
@@
[…]
@@
[…]
@@
Lemma 6.4.
Let $J_{n}$, $K_{n}$, $r_{n}$ be sequences of positive random variables such that $r_{n}$ converges almost surely to 0 and
@@
where $\widetilde{\gamma}_{n}=c_{\widetilde{\gamma}}n^{\widetilde{\gamma}}$ with $11/2<\widetilde{\gamma}<$ 1 and $c_{\widetilde{\gamma}}>0$. In addition, it is assumed that
@@
By applying Proposition A.4 in Godichon-Baggioni et al., (2023), we obtain:

### v3

@@ Appendix B Useful lemmas
Proof of Lemma B.2.
Note that the case where $E_{n}=0$ is exactly the case of Lemma B.1. Therefore, we are going to study the case where $E_{n}\neq 0$. We define $\alpha_{n}:=\prod_{k=0}^{n}(1+E_{k})$. Note that since $\sum_{n=0}^{\infty}E_{n}$ converges almost surely, $\alpha_{n}$ converges almost surely to a finite random variable $\alpha_{\infty}$. Moreover, noting
@@
$\mathbb{E}\left[V_{n+1}^{\prime}|\mathcal{F}_{n}\right]\leq V_{n}^{\prime}+B_{n}^{\prime}-D_{n}^{\prime}.$
@@
According to Lemma B.1, we have $V^{\prime}_{n}=o(a_{n})$ a.s., and therefore $V_{n}=o(a_{n})$ a.s. Furthermore,
@@
    $\left\|R_{n}\right\|_{F}=o\left(v_{n}\right)\quad a.s\quad\quad\text{where}\quad\quad v_{n}=\frac{(\ln n)^{a}}{n^{b}},$
@@
    $\beta_{n,k}A=\prod_{j=k+1}^{n}\left(I_{q}-\gamma_{j}\Gamma\right)A\prod_{j=k+1}^{n}\left(I_{q}-\gamma_{j}\Gamma\right)\quad\text{and}\quad\beta_{n,n}A=A,$
@@
Lemma B.4.
Let $J_{n}$, $K_{n}$, $r_{n}$ be sequences of positive random variables and $c$ be a positive constant such that $r_{n}$ converges almost surely to 00 and
@@
where $\widetilde{\gamma}_{n}=c_{\widetilde{\gamma}}n^{-\widetilde{\gamma}}$ with $11/2<\widetilde{\gamma}<$ 1 and $c_{\widetilde{\gamma}}>0$. In addition, it is assumed that
@@
By applying Proposition B.4 in Godichon-Baggioni et al., (2023), we obtain: